# On the Stability of Graph Convolutional Neural Networks: A Probabilistic Perspective

### v3

@@ 6 Conclusion
In this paper, we propose a novel probabilistic framework for analyzing the embedding stability of GCNNs under edge perturbations. Unlike prior worst-case analyses, our formulation enables a novel distribution-aware understanding of how graph structure and signal correlation jointly influence model sensitivity. This insight allows us to provide a structural interpretation of the importance of perturbation, and motivates the design of Prob-PGD, a new adversarial attack method shown to be effective through extensive experiments. While focused on inference-time stability, our study naturally supports broader robustness studies in graph machine learning, including adversarial training [32] by crafting adversarial examples. We leave such studies for future work.
@@ A.1 Summary of notations
${X}$ Random graph signal (node feature)
@@ A.2 Concept definitions
Table 3: Common graph filters from [39, 37] and their perturbations
@@
GIN [^3] We modify the standard GIN architecture by replacing the MLP in each layer with a single-layer perceptron, resulting in a simplified convolutional variant of GIN. [50] $(1+\epsilon^{(k)})\mathbf{I}+\mathbf{A}$ $\mathbf{A}-\mathbf{A}_{p}$ max degree
SGC($k$[49] $\left(\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}\right)^{k}$ $\left(\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}\right)^{k}-\left(\tilde{\mathbf{D}}_{p}^{-1/2}\tilde{\mathbf{A}}_{p}\tilde{\mathbf{D}}_{p}^{-1/2}\right)^{k}$ 1
Table 4: GCNNs and their associated graph filters.
@@ B.1 Proof of Theorem 1
We start with deriving the expression (4) for the expected embedding perturbation. Consider a random graph signal ${X}$ in $\mathbb{R}^{n}$, its $i$-th entry is a random variable, which we denote as $X_{i}$. Given a deterministic graph filter perturbation $\mathbf{E}_{g}$, the perturbation on the output embedding is $\mathbf{E}_{g}X$, which is also a random vector, and its squared Frobenius norm (Euclidean norm) can be written as
@@ B.3 Proof of Corollary 2
Under edge perturbations, the induced output embedding perturbation of a single-layered GCNN is more complicated than outputs of graph filters due to the existence of the non-linear activation function and the learned weight matrix $\mathbf{\Theta}^{(1)}\in\mathbb{R}^{1\times d}$.
@@
In (11), the inequality denotes entrywise less than, and ||$\cdot$|| ⋅ | denotes entrywise absolute value on the matrix. The inequality  (11) follows from the Lipschitz continuity of the activation function $\sigma(\cdot)$.
@@ B.5 Proof of Proposition 1
=$=\sum_{\{u,v\}\in\mathcal{P}}\mathbb{I}_{uv}^{2}+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}\left(\mathbb{I}_{uv}\mathbb{I}_{uv^{\prime}}+\mathbb{I}_{uv^{\prime}}\mathbb{I}_{uv}\right)+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{v^{\prime}u}\left(\mathbb{I}_{uv}\mathbb{I}_{v^{\prime}u}+\mathbb{I}_{v^{\prime}u}\mathbb{I}_{uv}\right)$
=$=\sum_{\{u,v\}\in\mathcal{P}}\mathbb{I}_{uu}+\mathbb{I}_{vv}+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}(\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v})+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}(\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v}).$ (28)
@@
=$=\sum_{\{u,v\}\in\mathcal{P}}\langle\mathbb{I}_{uu}+\mathbb{I}_{vv},\mathbf{K}\rangle+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}\langle\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v},\mathbf{K}\rangle+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}\langle\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v},\mathbf{K}\rangle$ (29)
=$=\sum_{\{u,v\}\in\mathcal{P}}\mathbf{K}_{uu}+\mathbf{K}_{vv}+2\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}\mathbf{K}_{vv^{\prime}}+2\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}\mathbf{K}_{vv^{\prime}}.$ (30)
@@
To start, we introduce the oriented edge indicator vector $\mathbf{b}_{uv}\in\{1,-1,0\}^{n}$, which only has ±$\pm 1$ on entries that correspond to two vertices $u$ and $v$ connected by an edge and zero elsewhere. For a directed graph,
$\mathbf{b}_{uv}(i)=\begin{cases}1&\text{if the edge points to $i$}\\ -1&\text{if the edge leaves $i$}\\ 0&\text{o.w.}\end{cases}$
@@
Here $\mathcal{P}$ denote the collection of ordered vertex pairs that are perturbed. For $\{u,v\}\in\mathcal{P}$, we denote $\sigma_{uv}=1$ if an edge is added to connect $u$ and $v$ and $\sigma_{uv}=-1$ if the edge between $u$ and $v$ is deleted. Then, we can further derive
@@
=$=\sum_{\{u,v\}\in\mathcal{P}}\mathbf{b}_{uv}\mathbf{b}_{uv}^{T}\mathbf{b}_{uv}\mathbf{b}_{uv}^{T}+\sum_{\begin{subarray}{c}\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}\end{subarray}}\sigma_{uv}\sigma_{uv^{\prime}}\mathbf{b}_{uv}\mathbf{b}_{uv}^{T}\mathbf{b}_{uv^{\prime}}\mathbf{b}_{uv^{\prime}}^{T}+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{v^{\prime}u}\mathbf{b}_{uv}\mathbf{b}_{uv}^{T}\mathbf{b}_{v^{\prime}u}\mathbf{b}_{v^{\prime}u}^{T}$
$=2\sum_{\{u,v\}\in\mathcal{P}}(\mathbb{I}_{uu}+\mathbb{I}_{vv}-\mathbb{I}_{uv}-\mathbb{I}_{vu})+\sum_{\begin{subarray}{c}\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}\end{subarray}}\sigma_{uv}\sigma_{uv^{\prime}}(2\mathbb{I}_{uu}+\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v}-\mathbb{I}_{uv^{\prime}}-\mathbb{I}_{v^{\prime}u}-\mathbb{I}_{vu}-\mathbb{I}_{uv})$
+$\;\ +\sum_{\{u,v\},(v^{\prime},u)\in\mathcal{P}}\sigma_{uv}\sigma_{v^{\prime}u}(2\mathbb{I}_{uu}+\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v}-\mathbb{I}_{uv^{\prime}}-\mathbb{I}_{v^{\prime}u}-\mathbb{I}_{vu}-\mathbb{I}_{uv})$ (32)
@@
$=2\sum_{\{u,v\}\in\mathcal{P}}(\mathbf{K}_{uu}+\mathbf{K}_{vv}-\mathbf{K}_{uv}-\mathbf{K}_{vu})+2\sum_{\begin{subarray}{c}\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}\end{subarray}}\sigma_{uv}\sigma_{uv^{\prime}}(\mathbf{K}_{uu}+\mathbf{K}_{vv^{\prime}}-\mathbf{K}_{uv^{\prime}}-\mathbf{K}_{vu})$ (33)
@@ C.1 Algorithm: Prob-PGD
Input : graph $\mathcal{G}$, second-moment matrix $\mathbf{K}$, edge perturbation budget $m$, maximum number of iterations $N$, learning rate $\alpha$, tolerance $\tau$
@@ C.2 Experiment setup
(8) Cora. The Cora dataset contains 2,708 scientific publications (vertices) categorized into 7 classes. These documents are connected by 5,429 citation links (edges), forming a citation network. Each document is described by a 1,433-dimensional sparse bag-of-words feature vector, where each dimension corresponds to a unique word from the corpus vocabulary.
@@ C.4 Comparison with the average-case metric for graph filters
Consider a graph signal ${X}$ sampled from the unit sphere in $\mathbb{R}^{n}$. Then the second-moment matrix $\mathbf{K}$ is positive semi-definite with $\mathrm{Tr}(\mathbf{K})=1$. Furthermore, the expected embedding perturbation satisfies
@@
Case 2. We consider the case where we have no prior information about the graph signal ${X}$ and simply assume that it is sampled uniformly at random from the unit sphere $\mathbb{S}^{n-1}$. Then we have that the second-moment matrix $\mathbf{K}=\frac{1}{n}\mathbf{I}$ [3].
@@ C.6 Experiments on message passing neural networks
MUTAG Unpert. 82.89$\pm$4.60 82.11$\pm$5.49 - -
Random 70.53$\pm$4.82 70.53$\pm$5.37 28.62$\pm$2.34 28.43$\pm$3.79
Wst-PGD 76.32$\pm$8.15 76.32$\pm$9.04 37.47$\pm$9.47 37.37$\pm$11.16
Prob-PGD 76.58$\pm$8.02 77.11$\pm$7.81 96.26$\pm$11.39 93.47$\pm$10.32
ENZYMES Unpert. 58.92$\pm$3.59 50.08$\pm$5.30 - -
Random 48.25$\pm$3.75 41.83$\pm$4.15 92.44$\pm$12.56 28.35$\pm$5.16
Wst-PGD 39.33$\pm$4.83 41.00$\pm$3.14 295.62$\pm$35.02 74.06$\pm$12.83
Prob-PGD 37.00$\pm$3.25 35.58$\pm$3.27 345.54$\pm$49.35 83.59$\pm$20.24
Cora Unpert. 73.76$\pm$1.95 70.72$\pm$2.30 - -
Random 69.73$\pm$2.63 66.95$\pm$3.16 1129$\pm$344 1498$\pm$1127
Wst-PGD 72.55$\pm$2.25 69.19$\pm$2.20 16529$\pm$7638 11105$\pm$6508
Prob-PGD 57.95$\pm$3.35 54.09$\pm$5.03 40343$\pm$23605 25784$\pm$13428

### v4

@@ 6 Conclusion
In this paper, we propose a probabilistic framework for analyzing the embedding stability of GCNNs under edge perturbations. Unlike prior worst-case analyses, our formulation naturally enables a distribution-aware understanding of how graph structure and signal correlation jointly influence model sensitivity. This insight allows us to provide a structural interpretation of the importance of perturbation, and motivates the design of Prob-PGD, a new adversarial attack method shown to be effective through extensive experiments. While focused on inference-time stability, our study naturally supports broader robustness studies in graph machine learning, including adversarial training [33] by crafting adversarial examples. We leave such studies for future work.
Acknowledgment
NZ acknowledges support from the Engineering and Physical Sciences Research Council (EPSRC) and IBM. XD acknowledges support from the Oxford-Man Institute of Quantitative Finance and EPSRC No. EP/T023333/1.
@@ A.1 Summary of notations
X{X} Random graph signal (node feature)
@@ A.2 Concept definitions
Table 3: Common graph filters from [42, 40] and their perturbations
@@
Table 4: GCNNs and their associated graph filters.
@@
GIN [53] [^3] We modify the standard GIN architecture by replacing the MLP in each layer with a single-layer perceptron, resulting in a simplified convolutional variant of GIN. $(1+\epsilon^{(k)})\mathbf{I}+\mathbf{A}$ $\mathbf{A}-\mathbf{A}_{p}$ max degree
SGC(kk[52] $\left(\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}\right)^{k}$ $\left(\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}\right)^{k}-\left(\tilde{\mathbf{D}}_{p}^{-1/2}\tilde{\mathbf{A}}_{p}\tilde{\mathbf{D}}_{p}^{-1/2}\right)^{k}$ 1
@@ B.1 Proof of Theorem 1
We start with deriving the expression (4) for the expected embedding perturbation. Consider a random graph signal X{X} in $\mathbb{R}^{n}$, its $i$-th entry is a random variable, which we denote as $X_{i}$. Given a deterministic graph filter perturbation $\mathbf{E}_{g}$, the perturbation on the output embedding is $\mathbf{E}_{g}X$, which is also a random vector, and its squared Frobenius norm (Euclidean norm) can be written as
@@ B.3 Proof of Corollary 2
Under edge perturbations, the induced output embedding perturbation of a single-layered GCNN is more complicated than outputs of graph filters due to the existence of the non-linear activation function and the learned weight matrix $\mathbf{\Theta}^{(1)}$.
@@
In (12), the inequality denotes entrywise less than, and $|\cdot|$ denotes entrywise absolute value on the matrix. The inequality  (12) follows from the Lipschitz continuity of the activation function $\sigma(\cdot)$.
@@ B.5 Proof of Proposition 1
$=\sum_{\{u,v\}\in\mathcal{P}}\mathbb{I}_{uv}^{2}+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}\left(\mathbb{I}_{uv}\mathbb{I}_{uv^{\prime}}+\mathbb{I}_{uv^{\prime}}\mathbb{I}_{uv}\right)+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{v^{\prime}u}\left(\mathbb{I}_{uv}\mathbb{I}_{v^{\prime}u}+\mathbb{I}_{v^{\prime}u}\mathbb{I}_{uv}\right)$
$=\sum_{\{u,v\}\in\mathcal{P}}\mathbb{I}_{uu}+\mathbb{I}_{vv}+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}(\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v})+\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\sigma_{uv}\sigma_{uv^{\prime}}(\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v}).$ (29)
@@
$=\sum_{\{u,v\}\in\mathcal{P}}\langle\mathbb{I}_{uu}+\mathbb{I}_{vv},\mathbf{K}\rangle+\hskip-16.0pt\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\hskip-12.0pt\sigma_{uv}\sigma_{uv^{\prime}}\langle\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v},\mathbf{K}\rangle+\hskip-16.0pt\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\hskip-12.0pt\sigma_{uv}\sigma_{uv^{\prime}}\langle\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v},\mathbf{K}\rangle$ (30)
$=\sum_{\{u,v\}\in\mathcal{P}}\mathbf{K}_{uu}+\mathbf{K}_{vv}+2\hskip-16.0pt\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\hskip-12.0pt\sigma_{uv}\sigma_{uv^{\prime}}\mathbf{K}_{vv^{\prime}}+2\hskip-16.0pt\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\hskip-12.0pt\sigma_{uv}\sigma_{uv^{\prime}}\mathbf{K}_{vv^{\prime}}.$ (31)
@@
To start, we introduce the oriented edge indicator vector $\mathbf{b}_{uv}\in\{1,-1,0\}^{n}$, which only has $\pm 1$ on entries that correspond to two vertices $u$ and $v$ connected by an edge and zero elsewhere. For a directed graph,
[…]
@@
Here $\mathcal{P}$ denote the collection of vertex pairs that are perturbed. For $\{u,v\}\in\mathcal{P}$, we denote $\sigma_{uv}=1$ if an edge is added to connect $u$ and $v$ and $\sigma_{uv}=-1$ if the edge between $u$ and $v$ is deleted. Then, we can further derive
@@
$=\sum_{\{u,v\}\in\mathcal{P}}\mathbf{b}_{uv}\mathbf{b}_{uv}^{T}\mathbf{b}_{uv}\mathbf{b}_{uv}^{T}+\hskip-12.0pt\sum_{\begin{subarray}{c}\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}\end{subarray}}\hskip-12.0pt\sigma_{uv}\sigma_{uv^{\prime}}\mathbf{b}_{uv}\mathbf{b}_{uv}^{T}\mathbf{b}_{uv^{\prime}}\mathbf{b}_{uv^{\prime}}^{T}+\hskip-12.0pt\sum_{\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}}\hskip-12.0pt\sigma_{uv}\sigma_{v^{\prime}u}\mathbf{b}_{uv}\mathbf{b}_{uv}^{T}\mathbf{b}_{v^{\prime}u}\mathbf{b}_{v^{\prime}u}^{T}$
$=2\sum_{\{u,v\}\in\mathcal{P}}(\mathbb{I}_{uu}+\mathbb{I}_{vv}-\mathbb{I}_{uv}-\mathbb{I}_{vu})+\hskip-16.0pt\sum_{\begin{subarray}{c}\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}\end{subarray}}\hskip-16.0pt\sigma_{uv}\sigma_{uv^{\prime}}(2\mathbb{I}_{uu}+\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v}-\mathbb{I}_{uv^{\prime}}-\mathbb{I}_{v^{\prime}u}-\mathbb{I}_{vu}-\mathbb{I}_{uv})$
$\;\ +\sum_{\{u,v\},(v^{\prime},u)\in\mathcal{P}}\sigma_{uv}\sigma_{v^{\prime}u}(2\mathbb{I}_{uu}+\mathbb{I}_{vv^{\prime}}+\mathbb{I}_{v^{\prime}v}-\mathbb{I}_{uv^{\prime}}-\mathbb{I}_{v^{\prime}u}-\mathbb{I}_{vu}-\mathbb{I}_{uv})$ (33)
@@
$=2\sum_{\{u,v\}\in\mathcal{P}}(\mathbf{K}_{uu}+\mathbf{K}_{vv}-\mathbf{K}_{uv}-\mathbf{K}_{vu})+2\hskip-16.0pt\sum_{\begin{subarray}{c}\{u,v\},\{u,v^{\prime}\}\in\mathcal{P}\end{subarray}}\hskip-16.0pt\sigma_{uv}\sigma_{uv^{\prime}}(\mathbf{K}_{uu}+\mathbf{K}_{vv^{\prime}}-\mathbf{K}_{uv^{\prime}}-\mathbf{K}_{vu})$ (34)
@@ C.1 Algorithm: Prob-PGD
Input : graph $\mathcal{G}$, second-moment matrix $\mathbf{K}$, edge perturbation budget $m$, maximum number of iterations NN, learning rate $\alpha$, tolerance $\tau$
@@ C.2 Experiment setup
(8) Cora. The Cora dataset contains 2708 scientific publications (vertices) categorized into 7 classes. These documents are connected by 5429 citation links (edges), forming a citation network. Each document is described by a 1433-dimensional sparse bag-of-words feature vector, where each dimension corresponds to a unique word from the corpus vocabulary.
@@ C.4 Comparison with the average-case metric for graph filters
Consider a graph signal X{X} sampled from the unit sphere in $\mathbb{R}^{n}$. Then the second-moment matrix $\mathbf{K}$ is positive semi-definite with $\mathrm{Tr}(\mathbf{K})=1$. Furthermore, the expected embedding perturbation satisfies
@@
Case 2. We consider the case where we have no prior information about the graph signal X{X} and simply assume that it is sampled uniformly at random from the unit sphere $\mathbb{S}^{n-1}$. Then we have that the second-moment matrix $\mathbf{K}=\frac{1}{n}\mathbf{I}$ [3].
@@ C.6 Experiments on message passing neural networks
[…]
C.7 Additional experiments on heterophilic datasets
We further conduct empirical study on the following two benchmark heterophilic datasets.
Chameleon. The Chameleon dataset contains 2277 nodes representing Wikipedia articles, where edges denote hyperlinks between them. Each node is assigned one of five classes according to the average monthly traffic of the corresponding page. With a homophily score of 0.23, reflecting strong heterophilic properties. Following the standard split, we use 1092 nodes for training, 729 for validation, and 456 for testing.
Squirrel. The Squirrel dataset contains 5201 nodes, each representing a Wikipedia article, with edges indicating hyperlinks between pages. Nodes are categorized into five classes based on the average monthly traffic. The dataset has a homophily score of 0.22, reflecting strong heterophilic properties. We follow the standard data split, using 2496 nodes for training, 1664 for validation, and 1041 for testing.
With low homophily scores, these two datasets are challenging for traditional GNNs. Therefore, we adopt the H2GCN model [58], a graph convolutional network specifically designed for heterophilic graphs. Given a node feature matrix $\mathbf{X}$ and graph adjacency matrix $\mathbf{A}$, H2GCN constructs node embedding as
$\mathbf{H}=[\mathbf{X};\mathbf{A}\mathbf{X};\mathbf{A}^{2}\mathbf{X}],$
where the three components correspond to the original node features $\mathbf{X}$, the aggregated 1-hop features $\mathbf{A}\mathbf{X}$, and the aggregated 2-hop features $\mathbf{A}^{2}\mathbf{X}$, respectively. The concatenated embedding $\mathbf{H}$ are then passed through a multi-layer perceptron (MLP) for node classification. For each dataset, we pre-train 10 H2GCNs and report in Table 6 the mean and standard deviation of both prediction accuracy and embedding perturbation under different attack methods by perturbing 1000 edges.
Table 6: Results on heterophilic datasets with H2GCN.
Dataset Method Prediction Accuracy (%) Embedding Pert. $\|\cdot\|_{F}$
Chameleon Unperturbed $55.92\pm 1.47$
Random $54.23\pm 1.16$ $136.65\pm 4.27$
Wst-PGD $54.87\pm 1.51$ $3099.12\pm 0.00$
Prob-PGD $\mathbf{52.96\pm 1.38}$ $\mathbf{6392.17\pm 0.00}$
Squirrel Unperturbed $50.38\pm 0.39$
Random $\mathbf{47.91\pm 0.47}$ $578.92\pm 3.86$
Wst-PGD $48.48\pm 0.73$ $1127.21\pm 0.00$
Prob-PGD $48.28\pm 0.73$ $\mathbf{3908.26\pm 0.00}$
Across both heterophilic graphs, the Prob-PGD attack consistently induces the largest embedding perturbation compared to baselines. On the Squirrel dataset, despite producing the most substantial embedding disruption, Prob-PGD does not lead to the largest degradation in prediction accuracy. This outcome is not unexpected: Prob-PGD is a task-agnostic attack that seeks maximal perturbation in node embeddings rather than direct misclassification. Consequently, the relationship between embedding perturbation and classification accuracy is not always linear.
C.8 Scalability study on large-scale graphs
Our proposed algorithm Prob-PGD has demonstrated efficiency through intensive experiments on medium-sized graphs. The underlying framework, however, is general and can be adapted to large-scale graphs with millions of nodes. In particular, our attack formulation (8) is not restricted to PGD-based optimization and can naturally incorporate alternative search strategies such as greedy heuristics. These approaches iteratively perturb the edge that induces the largest local embedding change and can operate on local subgraphs or within restricted edge subsets. By narrowing the perturbation space, such localized search substantially reduces computational overhead while maintaining competitive attack strength, providing a practical balance between efficiency and optimality.
To provide a concrete example, we adopt a scalable greedy edge-deletion approximation of the original optimization in (8). Specifically, we restrict the perturbation search to existing edges, randomly sample a subset of candidate edges at each iteration, and greedily delete the one that maximizes the attack objective in (8). This stochastic greedy strategy leverages graph sparsity for efficient computation. We denote our approximation by Prob and apply the same approximation to the worst-case baseline, denoted as Wst. We conduct experiments using the adjacency graph filter on four large-scale OGB benchmarks [20]: ogbn-arxiv, ogbn-mag, ogbn-products, and pokec, where ogbn-mag is heterogeneous and pokec exhibits label heterophily. For each dataset we perform adversarial edge-deletion attacks that remove 200 existing edges. We report average embedding perturbation, peak memory usage, and running time in Table 7.
Table 7: Scalability experiments on large-scale graphs.
Dataset Nodes Edges Embedding Pert. $\|\cdot\|_{F}$ Memory (MB) Time (s)
Random Wst Prob (Ours)
ogbn-arxiv 169,343 1,166,243 4.10 14.96 25.29 390.6 537.7
ogbn-mag 1,939,743 21,111,007 2.59 8.53 15.75 7362.4 1603.0
ogbn-products 2,449,029 61,859,140 17.65 24.39 193.84 5833.3 2798.5
pokec 1,632,803 30,622,564 292.32 303.30 334.20 18040.3 1172.1